A toy gun is aimed horizontally at a target on a wall X meters away. The plastic bullet hits Y

meters below the target. What was the initial speed of the bullet in meters per second?

Assume that g = 10 m/s?


1
Expert's answer
2021-11-02T10:01:07-0400

The initial speed is


"v=X/t=\frac{X\sqrt g}{\sqrt{2Y}}=2.236\frac{X}{\sqrt Y}"